OpenStudy (anonymous):

WARNING(Super HARD) Problem: FInd two rectangles for which the perimeter of the first is three times that of the second and the area of the second is three times that of the first. What are the integer values for the length and width of both of these rectangles ?

OpenStudy (anonymous):

I warned you.

OpenStudy (calculusxy):

I think that for the first rectangle (for the perimeter), it should the length of 3 and the width of 3. For the second one, the length should be 3 and width12 (12*3=36). Perimeter=3+3+3+3=12 Area=36 12*3=36

OpenStudy (anonymous):

But how would the first rectangle's perimeter be bigger than the second's ?

OpenStudy (anonymous):

It's simply impossible, how can a rectangle have a larger perimiter yet a smaller area than another with a smaller perimiter :\?

OpenStudy (anonymous):

No its not in so give me a format and hin to follow so I can solve it.

OpenStudy (calculusxy):

Well from what I understood, it asked for the area to be triple the amount of the first rectangle's perimeter.

OpenStudy (anonymous):

any rectangle with a larger perm will have a larger area.....

OpenStudy (anonymous):

If you think about it this way it maybe will make more sense. Call the length x and the width y. 2x+2y = perimiter. x*y = area If x gets bigger both will increase and the same goes for y. It's impossible for one shape to have a larger area than another yet have a smaller perimiter following that logic :D

OpenStudy (anonymous):

a + b = 3cd c+ d = 3ab Still can't solve this problem and it is possible.

OpenStudy (anonymous):

Did I say an hour I meant week sorry.

OpenStudy (anonymous):

IOnly got 4 mins left until

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

This time I got time

OpenStudy (mathstudent55):

Small perimeter with large area. Large perimeter with small area.